OpenStudy (anonymous):

I need help with this problem. Can anyone help? Write an equation for the volume, in cubic inches, of a bin with a square base and a height of 2 inches more than 3 times the length of the base. Use the equation to answer the following questions: a. When is the volume zero? List all possible values of x, that is, the length of a side of the base. b. Is there a maximum volume? If so, where does it occur? c. Is there a minimum volume? If so, where does it occur?

OpenStudy (anonymous):

h= 3x + 2 V =x*x*h = (x^2(3x+2)) =3x^3+2x^2 a. V=0 when x=0; all positive x b.&c. find V '; locate V ' =0; determine if V'=0 is the location of max or min by looking at the values of V' on either side: eg. positive to the left, negative to the right of the extremum means a max.

OpenStudy (anonymous):

Thank you for your help. I think I understand it. I was thinking along those lines. You confirmed my attempt at the equation, and for a. I hadn't really gotten to b & c.